OpenStudy (anonymous):

f(x)=2x+1 and g(x)=x^2-7, find (fg)(x) a. 2x^3+x^2-14x-7 b. 2x^2-13 c. 3x^2+7x-6 d. 2x^3-x^2-7x-7

OpenStudy (anonymous):

b)

OpenStudy (anonymous):

f(x^2-7)=2(x^2-7)+1=2x^2-13